OpenStudy (anonymous):

How is the process of adding and subtracting radicals similar to that of simplifying expressions with variables?

OpenStudy (anonymous):

@dpaInc please help :(

OpenStudy (precal):

like terms

OpenStudy (anonymous):

The two main operations you can perform when adding and subtracting radicals are: simplifying radicals for example\[\sqrt{27}=\sqrt{3\times9}=\sqrt{3}\times\sqrt{9} = 3\sqrt{3}\]and, once you have done this, gathering similar terms, for example \[\sqrt{2}+3\sqrt{2}=(1+3)\sqrt{2}=4\sqrt{2}\]Simplification is similar to the process in algebra of removing common factors from an expression, and gathering similar terms is very similar to gathering or combining terms in algebra.

OpenStudy (precal):

in radicals sometimes you can create like terms

OpenStudy (anonymous):

simplifying variable expressions is basically combining like terms: 3x +2y - 7x = -4x + 2y that's as simplified you can get.... it's the same with radicals: \(\large 3\sqrt2 + 2\sqrt7 - 7\sqrt2 = -4\sqrt2 + 2\sqrt7 \) and the same as variable expressions, that's as simple as you can get this...

OpenStudy (precal):